OpenStudy (anonymous):

Need help!!! Not sure if in right section, but I need help balancing equations __Fe+ ___O2->__Fe2O3

OpenStudy (aaronq):

so this is basically trial and error. you need to get the same number of each atom on both sides. The complications arise with the inclusion of the the subscripts. in \(O_2\) (the 2) and in \(Fe_2O_3\) work like the distributive property in math: \(2*(2x)=4x\) in chem it's the same: \(3~O_2\) means that there are 3*2=6 O atoms \(2~Fe_2O_3\) means that there 2*2=4 Fe atoms and 2*3= 6 O atoms